OpenStudy (anonymous):

Ms. Levy paid $189 for a hotel room and she bought two concert tickets priced at $99 each. How much money did Ms. Levy spend altogether? A student answered that Ms. Levy spent $576. A. ($200 + $100) × 2 = t The student's answer was reasonable. B. $200 + (2 × $100) = t The student's answer was not reasonable. C. $200 + ($100 ÷ 2) = t The student's answer was not reasonable. D. ($200 × 2) + $100 = t The student's answer was not reasonable.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

help somone

OpenStudy (michaelac959):

b i believe but the tickets were 99 and the hotel room 189 so where are they getting 200 and 100? are they just rounding

OpenStudy (michaelac959):

because the answer is 387

OpenStudy (anonymous):

can you help me with a different question

OpenStudy (anonymous):

The Jackson family drove 490 miles each day for 9 days. They then flew 1,292 miles. How far did the Jackson family travel? A student answered 5,702 miles. A. 500 × 9 + 1,300 = The student's answer was reasonable. B. 500 ÷ 9 + 1,300 = The student's answer was not reasonable. C. 400 × 9 + 1,200 = The student's answer was not reasonable. D. 400 ÷ 9 + 1,200 = The student's answer was not reasonable.

OpenStudy (michaelac959):

it has to be a because the students answer was spot on 490*9+1202=5702

OpenStudy (anonymous):

can you help me again

OpenStudy (michaelac959):

sure what grade is this?

OpenStudy (anonymous):

de

OpenStudy (michaelac959):

?

OpenStudy (anonymous):

4 th

OpenStudy (anonymous):

A giant freshwater catfish weighed 644 pounds. Ethan caught a fish that weighed 10 pounds. How many times heavier was the giant freshwater catfish than Ethan’s fish? A. 650 – 10 B. 650 + 10 C. 650 × 10 D. 650 ÷ 10

OpenStudy (michaelac959):

oh ok im in highschool so i got you

OpenStudy (anonymous):

i dont understand it bcuz even with all the rules of math it doesnt come out as any of the answer choices

OpenStudy (anonymous):

but it would come out as closest to D

OpenStudy (anonymous):

whati sthe most posple answer

OpenStudy (anonymous):

bcuz when we try to solve it i come out with 64.4 *10 and 650 ÷ 10 is the closest one i can see that comes to it bcuz it comes out as 65

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

you sure bcuz i would choose the one that is closest to the answer that i can get

OpenStudy (anonymous):

it is ok

OpenStudy (anonymous):

bye